An all-equity firm is considering the following projects:
Project Beta IRR
W .65       10.0 %
X .90       10.5
Y 1.20       14.0
Z 1.80       17.0
The T-bill rate is 5 percent, and the expected return on the market is 12 percent.
Required:
(a) Which projects have a higher expected return than the firm’s 12 percent cost of capital?

Project W has a (Click to select)lowerhigher expected return, Project X has a (Click to select)lowerhigher expected return, Project Y has a (Click to select)lowerhigher expected return, and Project Z has a (Click to select)higherlower expected return.

(b) Which projects should be accepted?

Project W should be (Click to select)acceptedrejected, Project X should be (Click to select)acceptedrejected, Project Y should be (Click to select)acceptedrejected, and Project Z should be (Click to select)acceptedrejected.

(c)

Which projects will be incorrectly accepted/rejected or correctly accepted/rejected if the firm's overall cost of capital were used as a hurdle rate?

Project W would be (Click to select)correctly acceptedcorrectly rejectedincorrectly acceptedincorrectly rejected, Project X would be (Click to select)correctly rejectedincorrectly acceptedincorrectly rejectedcorrectly accepted, Project Y would be (Click to select)incorrectly acceptedcorrectly rejectedincorrectly rejectedcorrectly accepted, and Project Z would be (Click to select)incorrectly rejectedincorrectly acceptedcorrectly rejectedcorrectly accepted.

"show steps please"
Answer #1

Expected Return of W =Risk free rate+Beta*(Market Return-Risk free Rate)=5%+0.0.65*(12%-5%) =9.55%
Expected Return of X =Risk free rate+Beta*(Market Return-Risk free Rate)=5%+0.0.90*(12%-5%) =11.30%
Expected Return of Y =Risk free rate+Beta*(Market Return-Risk free Rate)=5%+1.20*(12%-5%) =13.40%
Expected Return of Z =Risk free rate+Beta*(Market Return-Risk free Rate)=5%+1.80*(12%-5%) =17.60%

a. If cost of capital is 12%. When Expected rate is higher than cost of capital it is higher or else lower
Project W has lower expected return ,​​​​​​​​​​​​​​Project X has lower expected return .​​​​​​​​​​​​​​Project Y has higher expected return .
​​​​​​​​​​​​​​Project Z has higher expected return .

b.If IRR is less than expected return it should be accepted or it else it should be rejected.
​​​​​​​Project W would be  rejected ,​​​​​​​​​​​​​​​​​​​​​Project X would be accepted  .​​​​​​​​​​​​​​Project Y would be rejected. .
​​​​​​​​​​​​​​Project Z would be accepted.

c. If IRR would be used as hurdle rate
Project W Correctly rejected.
Project X incorrectly accepted​​​​​​​
Project Y incorrectly rejected. ​​​​​​​
Project Z correctly accepted